OpenStudy (anonymous):

Tom is 4 years older than Jerry. Nine years ago, Tom was 5 times as old as Jerry was then. How old is each now

OpenStudy (anonymous):

Suppose Tom's age is :x Jerry' age be: y x = 4 + y ---------1 Nine years ago, tom was (x-9) and Jerry was(y-9), So, x - 9 = 5(y - 9) 4 + y - 9 = 5y - 45 4y = 40 Calculate y and then x..

OpenStudy (anonymous):

thx